Shift report 20000729 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sat Jul 29 2000 - 08:22:20 EDT

  • Next message: http account: "Shift report 20000729 08:00-16:00"

    ------------------------------------------------------------
    Date: Sat Jul 29 08:22:19 2000, Shift: 00:00-08:00
    Supervisor: BS, with: ZIS
    Log book#3 pages 107 - 107
    
    Shift Summary:
    ------------------------------------------------------------
    No beam at all, quench problems all night. A few injections
    just after 0700, but no ramps.
    
    It appears that the listservers are down - no listmail
    comes through.
    
    pii6 has a problem with its keyboard cable - it has to
    be juggled at the back of the machine to start
    working again.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sat Jul 29 2000 - 10:47:42 EDT